# Dyson Equation for Correlated Linearizations and Test Error of Random Features Regression

### v2

@@ Abstract
This paper develops some theory of the Dyson equation for correlated linearizations and uses it to solve a problem on asymptotic deterministic equivalent for the test error in random features regression. The theory developed for the correlated Dyson equation includes existence-uniqueness, spectral support bounds, and stability properties. This theory is new for constructing deterministic equivalents for pseudo-resolvents of a class of correlated linear pencils. In the application, this theory is used to give a deterministic equivalent of the test error in random features ridge regression, in a proportional scaling regime, wherein we have conditioned on both training and test datasets.
60B20,
47N20,
68T07,
Anisotropic global law,
empirical test error,
Gaussian equivalence,
linearization,
Dyson equation,
pseudo-resolvent,
@@ 1 Introduction
Contemporary artificial neural networks have found widespread applications in diverse domains. A notable trend in modern neural network design is the increasing size and complexity of these models. In practical applications, there is a prevalent use of highly overparametrized models. These overparametrized models exhibit exceptional capacity, showcasing the ability to perfectly fit the training data, even in scenarios where the labels are pure noise and nonetheless generalize well [65]. Despite the remarkable practical success of neural networks, a considerable gap exists between theoretical understanding and real-world performance in machine learning. Neural networks pose unique challenges for analysis due to two key factors. Firstly, the high dimensionality of these models often leads to behaviors that defy conventional statistical knowledge. Secondly, the presence of non-linear activation functions, which are known to enhance the expressive capacity of neural networks, further complicates analysis.
@@
In recent years, random matrix theory has been used to provide valuable insights into the behavior of neural networks and other machine learning models. A notable line of research at the confluence of random matrix theory and neural networks combines the linear pencil method with operator-valued free probability to analyze simple neural networks [52, 1, 2, 3, 59]. This approach has been particularly fruitful in obtaining asymptotic characterizations of the training and test errors under various conditions.
@@
Beyond the linear model, the random features model introduced in [55] stands out as one of the simplest models with significant expressive power. In contrast to the linear model, the random features model incorporates a non-linear activation function and can be overparametrized. Despite its simplicity, the random features model provides a mathematically tractable framework that proves instrumental in studying phenomena observed in real-life machine learning models, such as multiple descent [51, 3, 16] and implicit generalization [20, 42]. Extensive studies have delved into the training and test error of random features ridge regression in high dimensions [29, 51, 1, 3, 36, 52, 50, 57, 59, 41, 62]. Notably, the non-linear random features model is connected to a simpler linear Gaussian model through a universality phenomenon [51, 29, 31, 40, 56, 41].
@@
The concept of linearization, or linear pencil method, entails representing rational functions of random matrices as blocks of inverses of larger random matrices which depend linearly on its random matrix inputs. These linearizations possess simpler correlation structures, rendering them more amenable to certain types of analysis. Beyond their application in analyzing the training and test error of neural networks, linearizations have been used to study polynomials of random matrices on the global scale [12, 15, 33, 38, 39] and the local scale [26, 28, 13]. This exploration naturally leads to the study of pseudo-resolvents, or generalized resolvents. Pseudo-resolvents are inherently more challenging to study than resolvents due to the absence of a spectral parameter spanning the entire diagonal. Nonetheless, one effective approach to study pseudo-resolvent involves analyzing a fixed-point equation known as the Dyson equation for linearizations (DEL) [12, 13, 26, 28].
@@ 1.1 Main Contributions
One of the main contributions of this work is to show the existence and uniqueness of the deterministic equivalent constructed from the DEL Theorem 2.1. This extends existing theory on Dyson equations which focuses on the cases of resolvents of correlated matrices and on pseudoresolvents of block matrices whose blocks are i.i.d. (or similarly mean-field); see Section 3 for further discussion on how this relates to existing DEL theory.
Furthermore, we give an approach to showing the stability of the DEL, which we implement for the empirical test error of random features ridge regression. We note that there is no known simple test for stability of linearizations, even in the case of i.i.d. blocks (see the discussion in [26]), and hence this must be checked for each problem to which this method is applied. Within this framework, we derive an anisotropic global law for a class of pseudo-resolvents with general correlation structures.
We then apply this general theory to derive an asymptotically exact representation for the empirical test error of random features ridge regression, conditioning on both the training and test datasets. Specifically, we confirm [46, Conjecture 1] under the additional assumption that the norms of the training and test random features matrices are bounded in expectation. As a consequence, we establish a general Gaussian equivalence principle for the empirical test error of random feature ridge regression. This Gaussian equivalence holds regardless of the choice of activation functions and labels, and we provide empirical evidence demonstrating that it accurately characterizes the training and test errors in real-world datasets.
@@ 1.2 Notation
Given a complex matrix $M$, we denote its real transpose by $M^{T}$ and the conjugate transpose by $M^{\ast}$. When dealing with matrix sub-blocks, $M_{i,j}^{\ast}=(M_{i,j})^{\ast}$ denotes the conjugate transpose of the $(i,j)$ sub-block of $M$. $\mathrm{Hol}(\mathcal{D},\mathcal{Y})$ denotes the set of holomorphic functions from a domain $\mathcal{D}$ to a normed vector space $\mathcal{Y}$.
[…]
@@ 1.3 Organization
Section 2 introduces the main results and settings of this work. In particular, Section 2.1 establishes the existence and uniqueness of a solution to the Dyson equation for correlated linearizations, alongside some key properties of the solution, such as a Stieltjes transform representation, under broad assumptions. Sections 2.2 and 2.3 then leverages this solution to derive conditions under which an anisotropic global law holds for the pseudo-resolvent of a linearization, offering a sketch of the proof to clarify and motivate these conditions. In Section 2.4, we apply these findings to a random features ridge regression model and present a deterministic equivalent for its empirical test error, alongside a discussion of the theorem and its implications. We position our work within the existing literature in Section 3. The proofs of the main results are provided in Sections 4, 5 and 6, respectively. Section 7 concludes the paper. Finally, we present preliminary results in Appendix A and the proof of some intermediate lemmas in Appendix B.
@@ 2.1 Dyson Equation for Correlated Linearizations
$L=\begin{bmatrix}A&B^{T}\\ B&Q\end{bmatrix}\in\mathbb{R}^{\ell\times\ell}.$ (1)
Here, $A\in\mathbb{R}^{n\times n}$ is self-adjoint, $Q\in\mathbb{R}^{d\times d}$ is self-adjoint with both $Q$ and $\mathbb{E}Q$ non-singular, and $B\in\mathbb{R}^{d\times n}$ is arbitrary. Our primary focus is analyzing the high-dimensional behavior of the pseudo-resolvent $(L-z\Lambda)^{-1}$, where $\Lambda:=\mathrm{diag}\{I_{n\times n},0_{d\times d}\}$ and $z\in\mathbb{H}:=\{z\in\mathbb{C}\,:\,\Im[z]>0\}$ represents the upper half of the complex plane. Our framework relies on the Dyson equation for linearizations (DEL)
@@
$\mathcal{S}:M\in\mathbb{C}^{\ell\times\ell}\mapsto\begin{bmatrix}\mathcal{S}_{1,1}(M)&\mathbb{E}[(A-\mathbb{E}A)M_{1,1}(B-\mathbb{E}B)^{T}]\\ \mathbb{E}[(B-\mathbb{E}B)M_{1,1}(A-\mathbb{E}A)]&\mathbb{E}[(B-\mathbb{E}B)M_{1,1}(B-\mathbb{E}B)^{T}]\end{bmatrix}\in\mathbb{C}^{\ell\times\ell}$ (3)
is a symmetric positivity-preserving linear map encoding the correlation structure of the linearization. Here, by positivity-preserving, we mean that $\mathcal{S}$ maps positive semidefinite matrices to positive semidefinite matrices. On the other hand, by symmetric, we mean that $(\mathcal{S}(M))^{\ast}=\mathcal{S}(M^{\ast})$ for every $M\in\mathbb{C}^{\ell\times\ell}$. The sub-superoperator $\mathcal{S}_{1,1}$ also preserves positivity. Importantly, the expectation in the superoperator is taken only over the linearization, and not over the superoperator’s argument.
In order to ensure the existence of a unique solution to the Dyson equation for linearizations, we need to restrict (2) to a suitable set. Consequently, based on properties of the pseudo-resolvent $(L-z\Lambda)^{-1}$, [^1] Under the present setting, $Q$ is non-singular and the Schur complement of the lower-right $d\times d$ block of $L-z\Lambda$ is given by $A-B^{T}Q^{-1}B-zI_{n}$. Since its imaginary part is negative definite, the Schur complement is non-singular. Hence, we may apply a block inversion lemma to conclude that the matrix $L-z\Lambda$ is non-singular, and the pseudo-resolvent $(L-z\Lambda)^{-1}$ is well-defined for every $z\in\mathbb{H}$. we introduce the admissible set.
@@
$\mathcal{M}_{+}:=\mathrm{Hol}(\mathbb{H},\mathcal{A}_{+}),\quad\mathcal{A}_{+}:=\{W\in\mathbb{C}^{\ell\times\ell}:\Im[W]\succ 0\}\cap\mathcal{A}.$ (6)
Our initial key result naturally revolves around establishing the existence and uniqueness of a solution for (2). In traditional Dyson equation theory, wherein the spectral parameter spans the entire diagonal, the existence of a unique solution usually emerges comprehensively from [37, Theorem 2.1]. Indeed, we leverage this theorem precisely to establish the existence and uniqueness of a solution to (5). However, due to the absence of a spectral parameter spanning the entire diagonal in our case, demonstrating the existence of a solution to (2) is not trivial and requires careful analysis. Nonetheless, by leveraging the suitable properties of the admissible set and the surrogate regularized Dyson equation, we obtain the following existence and uniqueness result.
@@
$\int_{\mathbb{R}}\Omega(\mathrm{d}\lambda)=\begin{bmatrix}I_{n}&-\mathbb{E}B^{T}(\mathbb{E}Q)^{-1}\\ -(\mathbb{E}Q)^{-1}\mathbb{E}B&(\mathbb{E}Q)^{-1}\mathbb{E}[BB^{T}](\mathbb{E}Q)^{-1}.\end{bmatrix}.$
Theorem 2.1 is essentially a combination of lemmas 4.17 and 4.9, and we refer the reader to these lemmas for the proof. For the rest this paper, we will utilize the notation $M$ to represent the unique solution as ensured by Theorem 2.1, $M^{(\tau)}$ to represent the unique solution to the regularized DEL (5), and we will omit the explicit mention of $z$ when the context confines it to a fixed $z\in\mathbb{H}$.
@@ 2.2 Assumptions for a General Anisotropic Global Law
    Gaussian Design. There exists $\gamma\in\mathbb{N}$, $g\sim\mathcal{N}(0,I_{\gamma})$, and a measurable symmetric [^2] By symmetric, we mean that $\mathcal{C}(x)=\mathcal{C}(x)^{T}$ for every $x\in\mathbb{R}^{\gamma}$. map $\mathcal{C}:\mathbb{R}^{\gamma}\mapsto\mathbb{R}^{\ell\times\ell}$ such that $L\equiv L(g)=\mathcal{C}(g)+\mathbb{E}L$.
@@
Before proceeding, we briefly discuss the implications of conditions (C2), (C1) and (C3). The first condition, (C1), ensures that key quantities remain bounded as the dimension of the problem increases. In the application of our framework, we will derive an explicit formula for the superoperator $\mathcal{S}$ in terms of the linearization $L$. This condition is a weaker form of the flatness condition commonly encountered in the Dyson equation literature [24, 8, 10]. Consequently, we will adopt the term flatness to describe this property. Specifically, since the blocks of the superoperator $\mathcal{S}$ are often quadratic in the blocks of the linearization $L$, the boundedness of $\|\mathcal{S}\|$ can be established through the boundedness of the second moments of the linearization $L$. This holds in common settings of interest where $L$ is correctly scaled. The boundedness of $\|\mathbb{E}L\|$ is a natural condition, ensuring that the linearization does not have a large deterministic component. Additionally, the boundedness of $\mathbb{E}\|(L-z\Lambda)^{-1}\|^{2}$ is a technical requirement that ensures the regularized pseudo-resolvent approximates the actual pseudo-resolvent effectively. Using the block inversion lemma, we can express
$(L-z\Lambda)^{-1}=\begin{bmatrix}R&-RB^{T}Q^{-1}\\ -Q^{-1}BR&Q^{-1}+Q^{-1}BRB^{T}Q^{-1}\end{bmatrix},$
@@
where $R=(A-B^{T}Q^{-1}B-zI_{n})^{-1}$. Given that $\|R\|\leq(\Im[z])^{-1}$, the condition $\limsup_{\ell\to\infty}\allowbreak\mathbb{E}\|(L-z\Lambda)^{-1}\|^{2}<\infty$ holds whenever, for example, $\|Q^{-1}\|$ and $\|B\|$ are bounded in $L^{8}$. Overall, the validity of condition (C1) can be established through boundedness of moments of the linearization $L$.
@@
The second condition, (C2), is a stability condition that ensures that the solution to the regularized Dyson equation for linearizations is a good approximation for the solution to the Dyson equation for linearizations. Stability is central to the analysis of Dyson equations, often studied via the stability operator. Following the notation of [10], the stability operator is defined as $\mathcal{L}:X\in\mathbb{C}^{\ell\times\ell}\mapsto X-M\mathcal{S}(X)M$, where $M\equiv M(z)$ is the unique solution to a Dyson equation. The stability operator plays a crucial role in establishing the stability of the Dyson equation, as shown in [5, Lemma 4.10] and others [24, 26]. The stability operator also naturally arises in the proof of Theorem 2.1, where its invertibility enables the recursive determination of coefficients in the power series expansion of the solution. The connection between the stability operator and condition (C2) becomes evident when considering the derivative of $M^{(\tau)}(z)$ with respect to $i\tau$, yielding $\mathcal{L}(\partial_{i\tau}M(z))=(M(z))^{2}$. Since $M(z)$ is bounded in operator norm (as a consequence of the Stieltjes transform representation in Theorem 2.1), having an invertible stability operator with a bounded inverse implies condition (C2). This condition is satisfied within the Dyson equation framework [26, 12, 28], as indicated in references [26, Equation 4.11][12, Estimates 6.3.3], and [28, Equation A.25]. Generally, condition (C2) holds due to the dimension-independent representation of the DEL solution, constructed using free probability techniques. For example, [33, Lemma 5.4] demonstrates that such a representation exists when $L$ has the form $L=A_{0}\otimes I_{n}+\sum_{j=1}^{k}A_{j}\otimes X_{j}$, with $\{A_{j}\}_{j=0}^{k}$ being self-adjoint matrices and $\{X_{j}\}_{j=1}^{k}$ independent random matrices drawn from i.i.d. centered Gaussian distributions.
@@
We note that while conditions (C1) and (C2) are in some sense essential to the approach we take, (C3) is not. It could be replaced by an argument that shows the error term in the DEL is small (for example by means of leave-one-out type analysis). We also note that similar assumptions have appeared in random matrix theory analyses of neural networks [46, Assumption 1]. Furthermore (C3) is frequently satisfied for problem setups arising in the random matrix theory of neural networks.

### v4

@@ Abstract
This paper develops some theory of the Dyson equation for correlated linearizations and uses it to solve a problem on asymptotic deterministic equivalent for the test error in random features regression. The theory developed for the correlated Dyson equation includes existence-uniqueness, spectral support bounds, and stability properties. This theory is new for constructing deterministic equivalents for pseudo-resolvents of a class of linearizations with correlated entries. In the application, this theory is used to give a deterministic equivalent of the test error in random features ridge regression, in a proportional scaling regime, wherein we have conditioned on both training and test datasets.
@@ 1 Introduction
Contemporary artificial neural networks have found widespread applications in diverse domains. A notable trend in modern neural network design is the increasing size and complexity of these models. In practical applications, there is a prevalent use of highly overparametrized models. These overparametrized models exhibit exceptional capacity, showcasing the ability to perfectly fit the training data, even in scenarios where the labels are pure noise and nonetheless generalize well [zhang_generalization]. Despite the remarkable practical success of neural networks, a considerable gap exists between theoretical understanding and real-world performance in machine learning. Neural networks pose unique challenges for analysis due to two key factors. Firstly, the high dimensionality of these models often leads to behaviors that defy conventional statistical knowledge. Secondly, the presence of non-linear activation functions, which are known to enhance the expressive capacity of neural networks, further complicates analysis.
@@
In recent years, random matrix theory has been used to provide valuable insights into the behavior of neural networks and other machine learning models. A notable line of research at the confluence of random matrix theory and neural networks combines linearizations with operator-valued free probability to analyze simple neural networks [mel2022anisotropic, Adlam2019ARM, tripledescent, adlam_doubledescent, tripuraneni2021overparameterization]. This approach has been particularly fruitful in obtaining asymptotic characterizations of the training and test errors under various conditions.
@@
Beyond the linear model, the random features model introduced in [Rahimi_RF] stands out as one of the simplest models with significant expressive power. In contrast to the linear model, the random features model incorporates a non-linear activation function and can be overparametrized. Despite its simplicity, the random features model provides a mathematically tractable framework that proves instrumental in studying phenomena observed in real-life machine learning models, such as multiple descent [montanari_RF, adlam_doubledescent, double_descent] and implicit generalization [chouard2022quantitative, jacot_RF]. Extensive studies have delved into the training and test error of random features ridge regression in high dimensions [pmlr-v119-gerace20a, montanari_RF, Adlam2019ARM, adlam_doubledescent, hastie_22, mel2022anisotropic, MEI20223, schröder2024asymptotics, tripuraneni2021overparameterization, hu2024asymptotics, wang2023overparameterized]. Notably, the non-linear random features model is connected to a simpler linear Gaussian model through a universality phenomenon [montanari_RF, pmlr-v119-gerace20a, goldt2022gaussian, hong_lu_GET, schroder2023deterministic, hu2024asymptotics].
@@
The concept of linearization—which also goes under the name “realization” in the control community [HELTON20181] and “linear pencil method” in operator theory [mel2022anisotropic, Adlam2019ARM, tripledescent, adlam_doubledescent, tripuraneni2021overparameterization]—entails representing rational functions of random matrices as blocks of inverses of larger random matrices which depend linearly on its random matrix inputs. These linearizations possess simpler correlation structures, rendering them more amenable to certain types of analysis. Beyond their application in analyzing the training and test error of neural networks, linearizations have been used to study polynomials of random matrices on the global scale [Anderson_2013, BelinschiMaiSpeicher+2017+21+53, haagerup_new_2005, HELTON20181, HELTON2006105] and the local scale [nemish_local_2020, fronk2023norm, anderson_anticommutator]. This exploration naturally leads to the study of pseudo-resolvents, or generalized resolvents. Pseudo-resolvents are inherently more challenging to study than resolvents due to the absence of a spectral parameter spanning the entire diagonal. Nonetheless, one effective approach to study pseudo-resolvent involves analyzing a fixed-point equation known as the Dyson equation for linearizations (DEL) [Anderson_2013, anderson_anticommutator, nemish_local_2020, fronk2023norm].
@@ 1.1 Main Contributions
One of the main contributions of this work is to show the existence and uniqueness of the deterministic equivalent constructed from the DEL in Theorem 2.1. This extends existing theory on Dyson equations which focuses on the cases of resolvents of correlated matrices and on pseudoresolvents of block matrices whose blocks are i.i.d. (or similarly mean-field); see Section 3 for further discussion on how this relates to existing DEL theory.
Furthermore, we give an approach to showing the stability of the DEL, which we implement for the empirical test error of random features ridge regression. We note that there is no known simple test for stability of linearizations, even in the case of i.i.d. blocks (see the discussion in [nemish_local_2020]), and hence this must be checked for each problem to which this method is applied. Within this framework, we derive an anisotropic global law for a class of pseudo-resolvents with general correlation structures.
We then apply this general theory to derive an asymptotically exact representation for the empirical test error of random features ridge regression, conditioning on both the training and test datasets. Specifically, we confirm [louart_random-matrix-approach, Conjecture 1] under the additional assumption that the norms of the training and test random features matrices are bounded in expectation. As a consequence, we establish a general Gaussian equivalence principle for the empirical test error of random feature ridge regression. This Gaussian equivalence holds regardless of the choice of activation functions and labels, and we provide empirical evidence demonstrating that it accurately characterizes the training and test errors in real-world datasets.
@@ 1.2 Notation
Given a complex matrix $M$, we denote its real transpose by $M^{\top}$ and the conjugate transpose by $M^{\ast}$. When dealing with matrix sub-blocks, $M_{i,j}^{\ast}=(M_{i,j})^{\ast}$ denotes the conjugate transpose of the $(i,j)$ sub-block of $M$. If $\mathcal{D}$ is a domain in a complex normed vector space $\mathcal{X}$ and $\mathcal{Y}_{0}$ is a subset of a complex normed vector space $\mathcal{Y}$, we denote by $\mathrm{Hol}(\mathcal{D},\mathcal{Y}_{0})$ the set of holomorphic functions $f:\mathcal{D}\to\mathcal{Y}$ with $f(\mathcal{D})\subseteq\mathcal{Y}_{0}$.
We use $\|\cdot\|$ to represent the standard Euclidean norm when applied to a vector and the spectral norm when applied to either matrices or complex matrix-valued functions. Additionally, we employ $\|\cdot\|_{F}$ to denote the Frobenius norm and $\|\cdot\|_{\ast}$ for the nuclear norm. The trace of a matrix is denoted by $\mathrm{tr}$.
@@ 1.3 Organization
Section 2 introduces the main results and settings of this work. In particular, Section 2.1 establishes the existence and uniqueness of a solution to the Dyson equation for correlated linearizations, alongside some key properties of the solution, such as a Stieltjes transform representation, under broad assumptions. Sections 2.2 and 2.3 then leverages this solution to derive conditions under which an anisotropic global law holds for the pseudo-resolvent of a linearization, offering a sketch of the proof to clarify and motivate these conditions. In Section 2.4, we apply these findings to a random features ridge regression model and present a deterministic equivalent for its empirical test error, alongside a discussion of the theorem and its implications. We position our work within the existing literature in Section 3. The proofs of the main results are provided in Sections 4, 5 and 6, respectively. Section 7 concludes the paper. Finally, we present preliminary results in A and the proof of some intermediate lemmas in B.
@@ 2.1 Dyson Equation for Correlated Linearizations
$L=\begin{bmatrix}A&B^{\top}\\ B&Q\end{bmatrix}\in\mathbb{R}^{\ell\times\ell}$. (1)
Here, $A\in\mathbb{R}^{n\times n}$ is self-adjoint, $Q\in\mathbb{R}^{d\times d}$ is self-adjoint with both $Q$ and $\mathbb{E}Q$ non-singular, and $B\in\mathbb{R}^{d\times n}$ is arbitrary. Our primary focus is analyzing the high-dimensional behavior of the pseudo-resolvent $(L-z\Lambda)^{-1}$, where $\Lambda:=\mathrm{diag}\{I_{n},0_{d\times d}\}$ and $z\in\mathbb{H}:=\{z\in\mathbb{C}\,:\,\Im[z]>0\}$ represents the upper half of the complex plane. Our framework relies on the Dyson equation for linearizations (DEL)
@@
$\mathcal{S}:M\in\mathbb{C}^{\ell\times\ell}\mapsto\begin{bmatrix}\mathcal{S}_{1,1}(M)&\mathbb{E}[(A-\mathbb{E}A)M_{1,1}(B-\mathbb{E}B)^{\top}]\\ \mathbb{E}[(B-\mathbb{E}B)M_{1,1}(A-\mathbb{E}A)]&\mathbb{E}[(B-\mathbb{E}B)M_{1,1}(B-\mathbb{E}B)^{\top}]\end{bmatrix}\in\mathbb{C}^{\ell\times\ell}$ (3)
is a positivity-preserving linear map encoding the correlation structure of the linearization. Here, by positivity-preserving, we mean that $\mathcal{S}$ maps positive semidefinite matrices to positive semidefinite matrices. The sub-superoperator $\mathcal{S}_{1,1}$ also preserves positivity. Importantly, the expectation in the superoperator is taken only over the linearization, and not over the superoperator’s argument.
Concretely, the more natural superoperator to consider would be the covariance-induced one
$M\mapsto\mathbb{E}[(L-\mathbb{E}L)M(L-\mathbb{E}L)^{\top}].$
However, in this form the off-diagonal and lower-right blocks depend on the full matrix $M$. Since our analysis requires that these blocks depend only on $M_{1,1}$ in order to close the self-consistent equations, we adopt the reduced version in (3) which is obtained by discarding the terms involving $M_{1,2},M_{2,1},$ and $M_{2,2}$ in the off-diagonal and lower-right blocks of $\mathbb{E}[(L-\mathbb{E}L)M(L-\mathbb{E}L)^{\top}]$. This reduction is justified under mild assumptions (e.g., uncorrelated blocks or independent rows of $B$ with well-conditioned covariance matrix), in which the discarded terms are negligible. The error associated with this approximation is explicitly controlled by the norm $\|\tilde{\mathcal{S}}\|$ appearing below. Finally, the precise form of $\mathcal{S}_{1,1}$ is not essential for the theory, as long as it is positivity-preserving and provides an asymptotically accurate approximation to the upper-left block of the covariance-induced superoperator. As a useful mental model, one can think of $\mathcal{S}_{1,1}(M)\approx\mathbb{E}[(L-\mathbb{E}L)M(L-\mathbb{E}L)^{\top}]_{1,1}$.
In order to ensure the existence of a unique solution to the Dyson equation for linearizations, we need to restrict (2) to a suitable set. Consequently, based on properties of the pseudo-resolvent $(L-z\Lambda)^{-1}$, [^1] Under the present setting, $Q$ is non-singular and the Schur complement of the lower-right $d\times d$ block of $L-z\Lambda$ is given by $A-B^{\top}Q^{-1}B-zI_{n}$. Since its imaginary part is negative definite, the Schur complement is non-singular. Hence, we may apply a block matrix inversion lemma (see Lemma A.9) to conclude that the matrix $L-z\Lambda$ is non-singular, and the pseudo-resolvent $(L-z\Lambda)^{-1}$ is well-defined for every $z\in\mathbb{H}$. we introduce the admissible set.
@@
$\mathcal{M}_{+}:=\mathrm{Hol}(\mathbb{H},\mathcal{A}_{+}),\quad\mathcal{A}_{+}:=\{W\in\mathbb{C}^{\ell\times\ell}:\Im[W]\succ 0\}.$ (6)
Our initial key result naturally revolves around establishing the existence and uniqueness of a solution for (2). In traditional Dyson equation theory, wherein the spectral parameter spans the entire diagonal, the existence of a unique solution usually emerges comprehensively from [helton2007operatorvalued, Theorem 2.1]. Indeed, we leverage this theorem precisely to establish the existence and uniqueness of a solution to (5). However, due to the absence of a spectral parameter spanning the entire diagonal in our case, demonstrating the existence of a solution to (2) is not trivial and requires careful analysis. Nonetheless, by leveraging the suitable properties of the admissible set and the surrogate regularized Dyson equation, we obtain the following existence and uniqueness result.
@@
$\int_{\mathbb{R}}\Omega(\mathrm{d}\lambda)=\begin{bmatrix}I_{n}&-\mathbb{E}B^{\top}(\mathbb{E}Q)^{-1}\\ -(\mathbb{E}Q)^{-1}\mathbb{E}B&(\mathbb{E}Q)^{-1}\mathbb{E}[BB^{\top}](\mathbb{E}Q)^{-1}.\end{bmatrix}$.
Theorem 2.1 is essentially a combination of lemmas 4.12 and 4.7, and we refer the reader to these lemmas for the proof. For the rest of this paper, we will utilize the notation $M$ to represent the unique solution as ensured by Theorem 2.1, $M^{(\tau)}$ to represent the unique solution to the regularized DEL (5), and we will omit the explicit mention of $z$ when the context confines it to a fixed $z\in\mathbb{H}$.
@@ 2.2 Assumptions for a General Anisotropic Global Law
    Gaussian Design. There exists $\gamma\in\mathbb{N}$, $g\sim\mathcal{N}(0,I_{\gamma})$, and a measurable symmetric [^2] By symmetric, we mean that $\mathcal{C}(x)=\mathcal{C}(x)^{\top}$ for every $x\in\mathbb{R}^{\gamma}$. map $\mathcal{C}:\mathbb{R}^{\gamma}\mapsto\mathbb{R}^{\ell\times\ell}$ with $\mathbb{E}[\mathcal{C}(g)]=0$ such that $L\equiv L(g)=\mathcal{C}(g)+\mathbb{E}L$.
@@
Before proceeding, we briefly discuss the implications of conditions (C2), (C1) and (C3). The first condition, (C1), ensures that key quantities remain bounded as the dimension of the problem increases. In the application of our framework, we will derive an explicit formula for the superoperator $\mathcal{S}$ in terms of the linearization $L$. This condition is a weaker form of the flatness condition commonly encountered in the Dyson equation literature [erdos2019matrix, alt_thesis, alt_Kronecker]. Consequently, we will adopt the term flatness to describe this property. Specifically, since the blocks of the superoperator $\mathcal{S}$ are often quadratic in the blocks of the linearization $L$, the boundedness of $\|\mathcal{S}\|$ can be established through the boundedness of the second moments of the linearization $L$. This holds in common settings of interest where $L$ is correctly scaled. The boundedness of $\|\mathbb{E}L\|$ is a natural condition, ensuring that the linearization does not have a large deterministic component. Additionally, the boundedness of $\mathbb{E}\|(L-z\Lambda)^{-1}\|^{2}$ is a technical requirement that ensures the regularized pseudo-resolvent approximates the actual pseudo-resolvent effectively. Using Lemma A.9, we can express
$(L-z\Lambda)^{-1}=\begin{bmatrix}R&-RB^{\top}Q^{-1}\\ -Q^{-1}BR&Q^{-1}+Q^{-1}BRB^{\top}Q^{-1}\end{bmatrix}$,
@@
where $R=(A-B^{\top}Q^{-1}B-zI_{n})^{-1}$. Given that $\|R\|\leq(\Im[z])^{-1}$, the condition $\limsup_{\ell\to\infty}\allowbreak\mathbb{E}\|(L-z\Lambda)^{-1}\|^{2}<\infty$ holds whenever, for example, $\|Q^{-1}\|$ and $\|B\|$ are bounded in $L^{8}$. Overall, the validity of condition (C1) can be established through boundedness of moments of the linearization $L$.
@@
The second condition, (C2), is a stability condition that ensures that the solution to the regularized Dyson equation for linearizations is a good approximation for the solution to the Dyson equation for linearizations. Stability is central to the analysis of Dyson equations, often studied via the stability operator. Following the notation of [alt_Kronecker], the stability operator is defined as $\mathcal{L}:X\in\mathbb{C}^{\ell\times\ell}\mapsto X-M\mathcal{S}(X)M$, where $M\equiv M(z)$ is the unique solution to a Dyson equation. The stability operator plays a crucial role in establishing the stability of the Dyson equation, as shown in [ajanki_stability_2019, Lemma 4.10] and others [erdos2019matrix, nemish_local_2020]. The stability operator also naturally arises in the proof of Theorem 2.1, where its invertibility enables the recursive determination of coefficients in the power series expansion of the solution. The connection between the stability operator and condition (C2) becomes evident when considering the derivative of $M^{(\tau)}(z)$ with respect to $i\tau$, yielding $\mathcal{L}(\partial_{i\tau}M(z))=(M(z))^{2}$. Since $M(z)$ is bounded in operator norm (as a consequence of the Stieltjes transform representation in Theorem 2.1), having an invertible stability operator with a bounded inverse implies condition (C2). This condition is satisfied within the Dyson equation framework [nemish_local_2020, Anderson_2013, fronk2023norm], as indicated in references [nemish_local_2020, Equation 4.11][Anderson_2013, Estimates 6.3.3], and [fronk2023norm, Equation A.25]. Generally, condition (C2) holds due to the dimension-independent representation of the DEL solution, constructed using free probability techniques. For example, [haagerup_new_2005, Lemma 5.4] demonstrates that such a representation exists when $L$ has the form $L=A_{0}\otimes I_{n}+\sum_{j=1}^{k}A_{j}\otimes X_{j}$, with $\{A_{j}\}_{j=0}^{k}$ being self-adjoint matrices and $\{X_{j}\}_{j=1}^{k}$ independent random matrices drawn from i.i.d. centered Gaussian distributions.
@@
We note that while conditions (C1) and (C2) are in some sense essential to the approach we take, (C3) is not. It could be replaced by an argument that shows the error term in the DEL is small (for example by means of leave-one-out type analysis333“Leave-one-out” methods refer to standard techniques in random matrix theory where one decomposes the resolvent of a matrix into that of a closely related matrix obtained by removing a single row, column, or entry, together with an additive perturbation term in which the removed part is decoupled from the resolvent (see, e.g., the proof of the Marchenko-Pastur law in [bai2010spectral, Section 3.3.2], which is referred to as “leave-one-out” in [Couillet_Liao_2022]). This approach is related to Lindeberg-type replacement arguments, but differs in that one removes, rather than replaces, rows, columns, or entries.). We also note that similar assumptions have appeared in random matrix theory analyses of neural networks [louart_random-matrix-approach, Assumption 1]. Furthermore (C3) is frequently satisfied for problem setups arising in the random matrix theory of neural networks.